what is the x and y intercepts? x/a + y/b = 1
Answer by edjones(8007)

x/a + y/b = 1
y/b = 1 - x/a
y = b - xb/a
y=b when x=0, the y intercept.
.
y = b - xb/a
b - xb/a = 0
ab - xb = 0 multiply each side by a
a-x=0 divide b by each side.
-x=-a
x=a the x intercept when y=0.
.
Ed